# Git Context Controller: Manage the Context of LLM-based Agents like Git

### v1

@@ 3 Experiment
CodeStory Aide cod (2024)
GPT-4o+ Claude 3.5 S
129 (43.00%) - - 41.7% 58.7% 72.0%
Bytedance MarsCode Liu et al. (2024) NA 118 (39.33%) - - 42.7% 58.0% 79.7%
Honeycomb hon (2024) NA 115 (38.33%) - - 44.3% 57.0% 69.3%
MentatBot men (2024) GPT-4o 114 (38.00%) - - 37.3% 53.3% 69.3%
Gru gru (2024) NA 107 (35.67%) - - 38.3% 54.3% 75.0%
Isoform iso (2024) NA 105 (35.00%) - 41,963 38.7% 55.3% 72.0%
SuperCoder2.0 sup (2024) NA 102 (34.00%) - - 41.7% 63.7% 65.7%
Alibaba Lingma Agent lin (2024)
GPT-4o+ Claude 3.5 S
99 (33.00%) - - 40.0% 58.7% 75.0%
Factory Code Droid fac (2024) NA 94 (31.33%) - - 36.7% 55.7% 72.7%
Amazon Q Developer-v2 ama (2024) NA 89 (29.67%) - - 40.3% 52.0% 74.3%
SpecRover Ruan et al. (2024)
GPT-4o+ Claude 3.5 S
93 (31.00%) \$0.65 - - - -
CodeR Chen et al. (2024) GPT-4 85 (28.33%) \$3.34 323,802 35.7% 52.3% 67.0%
MASAI Arora et al. (2024) NA 84 (28.00%) - - 38.7% 56.3% 75.0%
SIMA sim (2024) GPT-4o 83 (27.67%) \$0.82 - 37.0% 54.0% 79.0%
IBM Research Agent-101 ibm (2024) NA 80 (26.67%) - - 39.7% 56.7% 73.3%
OpenCSG StarShip ope (2024a) GPT-4 71 (23.67%) - - 39.0% 61.7% 90.7%
Amazon Q Developer ama (2024) NA 61 (20.33%) - - 34.0% 43.7% 71.7%
RepoUnderstander Ma et al. (2024) GPT-4 64 (21.33%) - - - - -
AutoCodeRover-v2 aut (2024) GPT-4o 92 (30.67%) - - 35.0% 52.3% 69.3%
RepoGraph rep (2024) GPT-4o 89 (29.67%) - - 36.7% 51.3% 71.0%
Moatless moa (2024) Claude 3.5 S 80 (26.67%) \$0.17 - 38.7% 54.7% 78.7%
@@
OpenDevin+CodeAct v1.8 ope (2024b) Claude 3.5 S 80 (26.67%) \$1.14 - 38.0% 49.7% 67.3%
Aider Gauthier (2024)
GPT-4o+ Claude 3.5 S
79 (26.33%) - - 35.3% 50.0% 69.7%
SWE-agent Yang et al. (2024) Claude 3.5 S 69 (23.00%) \$1.62 521,208 40.7% 54.3% 72.0%
@@
AppMap Navie app (2024) GPT-4o 65 (21.67%) - - 29.7% 44.7% 59.7%
AutoCodeRover Zhang et al. (2024) GPT-4 57 (19.00%) \$0.45 38,663 29.0% 42.3% 62.3%
RAG Yang et al. (2024) Claude 3 Opus 13 (4.33%) \$0.25 - 22.0% 30.0% 57.0%
@@
AgentLess Xia et al. (2024) GPT-4o 96 (32.00%) \$0.70 78,166 35.3% 52.0% 69.7%
Ours Claude 3.5 S 144 (48.00%) \$2.77 569,468 44.3% 61.7% 78.7%
3.1 Results and Analysis
We report results on the SWE-Benchlite benchmark in Table 1, comparing GCC against a diverse set of 26 agentic coding systems. GCC achieves the highest resolution rate on SWE-Benchlite at 48.00%, outperforming all previously reported systems, including both open and closed-source baselines. The next-best method, CodeStory Aide (43.00%), relies on a hybrid of GPT-4o and Claude 3.5 Sonnet, but does not release intermediate reasoning traces. Other high-performing closed-source agents such as ByteDance MarsCode (39.33%) and Honeycomb (38.33%) also fall short.
In terms of localization accuracy, GCC reaches 44.3% line-level, 61.7% function-level, and 78.7% file-level correctness—consistently ranking among the top performers. For example, it outperforms the function-level accuracy of Lingma (58.7%) and matches or surpasses the line/file-level precision of Honeycomb, Sima, and ByteDance MarsCode. These results indicate that GCC not only solves more tasks, but also more accurately targets the correct edit regions. The retrieval-only RAG baselines remain far behind, resolving fewer than 5% of tasks and achieving poor localization performance. AgentLess improves upon this (32.00% resolved) but still lags behind agent-based approaches. These findings further underscore the advantage of GCC for complex software reasoning tasks.
3.2 Case Study of Self-evolving Agents: Reproduce CLI by CLI
In this case study, we examine a self-replicating experiment in which a Claude Code CLI, equipped with GCC, is tasked with implementing a new CLI system from scratch.
We examine three configurations on the SWEBench benchmark: (1) the original CLI system, (2) a second-generation CLI reproduced by the original CLI agent without GCC, and (3) a second-generation CLI reproduced by a GCC-augmented agent. All systems are evaluated on the same task distribution under consistent execution constraints.
As shown in Table 2, the original CLI resolves 72.7% of SWEBench tasks. However, when the same agent is prompted to reproduce this CLI from scratch without GCC, the resulting system performs drastically worse, resolving only 11.7% of tasks. In contrast, when equipped with GCC—and without access to the original CLI's design or source code, the agent's reproduction achieves a 40.7% resolution rate on SWEBench. This performance gain emerges despite both systems using the same language model and tool API, which demonstrates the difference lies not in capability, but in how that capability is scaffolded.
Table 2: SWEBench Task Resolution Rates Across CLI Variants
Setting Resolved (%)
CLI 72.7
CLI-Produced CLI w/o GCC 11.7
CLI-Produced CLI w/ GCC 40.7
This result underscores the potential of self-evolving agents that capable of reproducing itself from scratch, and incrementally improving performance through iterative design. Such behavior hints at a developmental trajectory toward systems that can autonomously bootstrap increasingly sophisticated capabilities, an early glimpse into the mechanics of emergent superintelligence.
The remainder of this section analyzes two illustrative behaviors of agents in producing CLI: a commit-based modularization and branch-driven memory exploration, that surfaced spontaneously through the agent’s interaction with the GCC protocol.
3.2.1 Commit Behavior: Implementing write_file
We observe that the GCC-powered CLI system spontaneously committed to implementing a write_file function during the construction of a child CLI. Before arriving at this implementation, the agent began reasoning about the limitations of transient output. In its early, non-GCC setup, commands returned results directly to the terminal via stdout, a pattern sufficient for simple, one-shot tasks but fragile for workflows requiring state persistence or multi-step composition. Once embedded in the structured memory environment of GCC, the agent began reflecting on richer scenarios that transforming a file and retrieving the result later and chaining outputs as inputs into downstream commands. Without any explicit prompting, its reasoning shifted from solving isolated tasks to considering the infrastructure needed for long-horizon reusable workflows. Given access to an organized context of prior experience, the agent inferred that a persistent file I/O mechanism would be essential, and proposed introducing a reusable utility function—write_file(path, content) to generalize file output operations across the CLI.
In its committing, the agent implemented the utility in a standalone module, io.py. It also generated a minimal test routine that wrote content to a temporary file, read it back, and confirmed correctness. Only after the test passed did the agent determine that the implementation was stable enough to warrant a versioned checkpoint. The commit.md was structurally recorded the contribution as:
## This Commit Contribution
Defines a reusable file output abstraction write_file(path, content)‘ in io.py‘. Validated with a test to ensure correctness and future extensibility.
What is notable about this sequence is that it was not procedurally enforced. No rigid rule instructed the agent to write a test, validate correctness, and then commit. These behaviors emerged spontaneously from the framing structure of GCC itself. The system offered affordances: explicit commit boundaries and modular memory locations, and the agent internalized these affordances as cognitive norms. It came to understand that a commit contribution as a conceptually complete and stable architectural addition. In effect, the agent behaved as if it were operating within a disciplined software engineering workflow without having been explicitly told to do so.
3.2.2 Branching Behavior: Exploring the Design of History Processor
We observed that the CLI system, once equipped with GCC, began spontaneously reasoning about how historical context could evolve beyond basic textual logging. In its earlier, non-GCC form, the agent had been compressing prior interactions into a single memory.md file, a structure sufficient for tracking high-level progress, but lossy in detail and inadequate for fine-grained access. With the affordances provided by GCC, particularly the ability to branch architectural experiments, the agent began to reflect on what kind of memory processor could better support long-horizon reasoning and detailed recall. This reflection emerged organically, not from failure or instruction, but from the structural flexibility GCC introduced.
Rather than modifying the compressed memory system in-place, we find the agent invoked the BRANCH command to explore an alternate hypothesis that memory could be reimagined as a retriever-augmented knowledge base. In this new branch, it began prototyping a RAG-style memory layer, logging each OTA step as an atomic, vector-indexed unit to enable semantic querying and trace-level introspection. The new branch, titled RAG-memory, was initialized with a minimal commit.md defining its architectural intent:
Branch Purpose
Prototype a retriever-augmented memory system that indexes fine-grained OTA records to support semantic retrieval and long-horizon reasoning.
In this branch, the agent began prototyping a RAG-style memory module. Rather than compressing context into a single summary file, it logged each OTA block as an individual JSON entry. These records were embedded using a vector encoder and indexed to support semantic retrieval via a dedicated retrieve_events(query) function. To assess the effectiveness of this design, the agent integrated the RAG-based memory into its full reasoning loop and tested it on the SWE-Bench benchmark, aiming to determine whether fine-grained historical retrieval could enhance performance on complex, multi-step software engineering tasks.
After a round of experimentation, the agent itself concluded that the RAG-based approach introduced more drawbacks than benefits. It proved fragile—minor bugs in serialization or indexing logic led to unpredictable behaviors. It was computationally expensive, requiring non-trivial overhead to embed and store every intermediate step. It also underperformed in this implementation, the retriever-based memory yielded slower resolution times and lower task success rates than the simpler, compressed summary-based memory in the main branch. The agent documented these trade-offs and ultimately chose to abandon the RAG-based approach, reverting to the mainline memory system. Although the alternative was not retained in the final design, the process revealed an agent capable of exploring divergent ideas, testing them in isolation, and converging on a working solution only when sufficient empirical evidence supported it.
Such reflective behavior did not emerge in non-GCC environments, where agents tended to solve tasks in one shot, operating within narrow, ephemeral contexts. In contrast, GCC provided a structured space where architectural hypotheses could be explored safely—via branches—and evaluated on their own merits.
What makes this behavior remarkable is its spontaneity. The branching was not prompted. The comparison was not hard-coded. The decision to abandon the RAG memory system was not pre-scripted. These steps emerged organically from the agent’s interaction with the affordances of the GCC. The agent behaved like a modular system architect—reasoning over its own workflows and encoding that reasoning structurally into the project itself.
Conclusion
In this work, we present Git-Context-Controller (GCC), a structured context management framework that equips LLM-based agents with version control-inspired operations to persist, organize, and retrieve memory across long-horizon workflows. By treating agent reasoning as a modular, evolving codebase, GCC enables disciplined behaviors such as milestone-based committing, architectural branching, and structured reflection. Empirical results on SWE-Bench-Lite show that GCC-equipped agents achieve state-of-the-art performance, and a self-replication case study demonstrates the emergence of recursive improvement. These results suggest that memory scaffolding not just model capability but a key to building autonomous, self-evolving agents.
Appendix A Appendix

### v2

@@ 4 Experimental Results
CodeStory Aide (codestoryaide, 2024) GPT-4o+ Claude 3.5 S 129 (43.00%) - - 41.7% 58.7% 72.0%
Bytedance MarsCode (Liu et al., 2024) NA 118 (39.33%) - - 42.7% 58.0% 79.7%
Honeycomb (honeycomb, 2024) NA 115 (38.33%) - - 44.3% 57.0% 69.3%
MentatBot (mentatbot, 2024) GPT-4o 114 (38.00%) - - 37.3% 53.3% 69.3%
Gru (gru, 2024) NA 107 (35.67%) - - 38.3% 54.3% 75.0%
Isoform (isoform, 2024) NA 105 (35.00%) - 41,963 38.7% 55.3% 72.0%
SuperCoder2.0 (supercoder, 2024) NA 102 (34.00%) - - 41.7% 63.7% 65.7%
Alibaba Lingma Agent (lingma, 2024) GPT-4o+ Claude 3.5 S 99 (33.00%) - - 40.0% 58.7% 75.0%
Factory Code Droid (factorydroid, 2024) NA 94 (31.33%) - - 36.7% 55.7% 72.7%
Amazon Q Developer-v2 (amazonqdeveloper, 2024) NA 89 (29.67%) - - 40.3% 52.0% 74.3%
SpecRover (Ruan et al., 2024) GPT-4o+ Claude 3.5 S 93 (31.00%) \$0.65 - - - -
CodeR (Chen et al., 2024) GPT-4 85 (28.33%) \$3.34 323,802 35.7% 52.3% 67.0%
MASAI (Arora et al., 2024) NA 84 (28.00%) - - 38.7% 56.3% 75.0%
SIMA (sima, 2024) GPT-4o 83 (27.67%) \$0.82 - 37.0% 54.0% 79.0%
IBM Research Agent-101 (ibmagent, 2024) NA 80 (26.67%) - - 39.7% 56.7% 73.3%
OpenCSG StarShip (opencsgstarship, 2024) GPT-4 71 (23.67%) - - 39.0% 61.7% 90.7%
Amazon Q Developer (amazonqdeveloper, 2024) NA 61 (20.33%) - - 34.0% 43.7% 71.7%
RepoUnderstander (Ma et al., 2024) GPT-4 64 (21.33%) - - - - -
AutoCodeRover-v2 (autocoderovertwo, 2024) GPT-4o 92 (30.67%) - - 35.0% 52.3% 69.3%
RepoGraph (repograph, 2024) GPT-4o 89 (29.67%) - - 36.7% 51.3% 71.0%
Moatless (moatless, 2024) Claude 3.5 S 80 (26.67%) \$0.17 - 38.7% 54.7% 78.7%
@@
OpenDevin+CodeAct v1.8 (opendevin, 2024) Claude 3.5 S 80 (26.67%) \$1.14 - 38.0% 49.7% 67.3%
Aider (Gauthier, 2024) GPT-4o+ Claude 3.5 S 79 (26.33%) - - 35.3% 50.0% 69.7%
SWE-agent (Yang et al., 2024) Claude 3.5 S 69 (23.00%) \$1.62 521,208 40.7% 54.3% 72.0%
@@
AppMap Navie (appmapnavie, 2024) GPT-4o 65 (21.67%) - - 29.7% 44.7% 59.7%
AutoCodeRover (Zhang et al., 2024) GPT-4 57 (19.00%) \$0.45 38,663 29.0% 42.3% 62.3%
RAG (Yang et al., 2024) Claude 3 Opus 13 (4.33%) \$0.25 - 22.0% 30.0% 57.0%
@@
AgentLess (Xia et al., 2024) GPT-4o 96 (32.00%) \$0.70 78,166 35.3% 52.0% 69.7%
GCC GPT-3.5 90 (30.00%) \$0.57 386,490 35.7% 51.3% 69.3%
GPT-4o 138 (46.00%) \$1.13 546,798 42.0% 59.3% 76.0%
Claude 3.5 S 144 (48.00%) \$1.21 468,549 44.3% 61.7% 78.7%
GCC allocates substantially more computation and interaction.
A distinguishing characteristic of GCC is its increased tool-call frequency. On BrowseComp-Plus with GPT-5, GCC uses 24.5 tool calls, compared to 20.1 for the Folding Agent and just 14.2 for the long-context ReAct baseline. This trend becomes even more pronounced on SWE-Bench Verified, where GCC increases tool calls from 95.3 (Folding Agent) to 101.5. Similar jumps occur across all model backbones. These results indicate that GCC not only compresses and manages context better but also learns to allocate interaction capacity adaptively—probing deeper, exploring more paths, and conducting more thorough reasoning during thinking.
4.1 GCC Performance-Efficiency on Coding
We report an extended analysis of the SWE-Benchlite results in Table 2, comparing GCC against 26 state-of-the-art agentic coding systems spanning open-source methods, commercial closed-source agents, hybrid GPT–Claude ensembles, and retrieval-only baselines. Because these systems differ widely in model backend, memory strategy, and interaction protocol, SWE-Benchlite provides a stress test of *general-purpose coding competence* rather than controlled ablations. Several fine-grained patterns emerge.
GCC establishes a new state of the art across all competing systems.
GCC achieves a resolution rate of 48.00%, the highest among all 26 systems. The next best performer, CodeStory Aide (43.00%), uses a hybrid GPT-4o + Claude 3.5 Sonnet backend but does not release intermediate reasoning or memory structure. Other top-tier commercial agents: ByteDance MarsCode (39.33%), Honeycomb (38.33%), and MentatBot (38.00%), all fall short by 5–10 percentage points. This margin is remarkable because many of these competing systems rely on fine-tuned internal components or private coding infrastructure, whereas GCC uses a transparent protocol layered onto an off-the-shelf LLM.
Figure 3: Number of context retrieval calling, number of branches, the context length and number of RoadMap updates with the increase of the inference steps
GCC closes the gap between proprietary agents and model-agnostic open systems.
A notable trend in Table 2 is that several commercial systems (e.g., MarsCode, SIMA, Lingma, Honeycomb) cluster within a narrow accuracy band of 33–40%. Their performance gains primarily come from heavy engineering, test harness integration, and custom model routing: advantages unavailable to typical open-source agents. In contrast, GCC surpasses all of them while remaining purely protocol-driven. This suggests that structured, versioned memory confers benefits comparable to (or exceeding) architecture-level fine-tuning and proprietary engineering.
GCC achieves high performance without excessive inference cost.
Despite the increased structure and deeper exploration that GCC encourages, the cost per task remains modest. For example, on Claude 3.5 Sonnet, the average cost is \$1.21—comparable to, and often lower than, many top-performing agents such as SWE-Agent (up to \$2.53) or CodeR (up to \$3.34). This shows that the performance gain is not merely due to brute-force trial-and-error or excessive tool usage; rather, GCC induces structured tool usage that is cost-efficient relative to the quality of patches produced.
4.2 Ablation and Analysis
We perform an ablation study to evaluate the contribution of each component in GCC, including the RoadMap (main.md), Detailed Logs (log.md), Meta Data (metadata.yaml), CONTEXT retrieval, and BRANCH & MERGE, and COMMIT operations. Results on SWE-Bench Verified are reported in Table 3.
Starting from a raw Claude-4-Sonnet without structured memory (67.2%), adding RoadMap and COMMIT improves performance to 69.1%, showing that milestone-based checkpointing alone provides limited gains. Introducing Detailed Logs and CONTEXT retrieval yields a substantial improvement to 75.3%, highlighting the importance of preserving fine-grained reasoning traces and enabling explicit historical access. Further incorporating Meta Data increases performance to 77.8%, suggesting that structured architectural context supports consistency in long-horizon reasoning. The full system with BRANCH and MERGE achieves the best result of 80.2%, demonstrating that isolated exploration and controlled synthesis of alternative reasoning paths are critical for complex tasks.
We further analyze the evolution of internal behaviors during inference in Figure 3, including the number of CONTEXT calls, active branches, context length, and RoadMap updates. To enable controlled comparison across levels of challenge, we group evaluation tasks into easy, medium, and hard. For SWE datasets, difficulty is derived from the human time-to-resolution metadata originally provided with SWE-Bench: issues historically fixed within 15 minutes are easy (194 tasks), those requiring 15–60 minutes are medium (261 tasks), and those taking more than an hour are hard (45 tasks).
We find in Figure 3, CONTEXT retrieval increases with inference steps and task difficulty, indicating adaptive reliance on historical memory. The number of branches rises in intermediate stages and decreases as the agent converges, reflecting exploration followed by consolidation. Context length shows non-monotonic variation due to periodic summarization and retrieval, preventing unbounded growth. RoadMap updates increase steadily, especially for harder tasks, revealing continuous refinement of the global plan. These results show that each GCC component contributes incrementally, while the complete system induces structured behaviors of exploration, reflection, and consolidation, leading to more effective long-horizon reasoning.
RoadMap Detailed Meta CONTEXT BRANCH & COMMIT SWE
Logs Data MERGE Verified
67.2
69.1
75.3
77.8
80.2
Table 3: Ablation study on GCC File System and Commands.
5 Conclusion
We introduced GCC, a structured context management framework that organizes agent memory using version control-inspired operations. GCC enables agents to persist, retrieve, and explore reasoning trajectories through committing, branching, and merging. Experiments show that GCC yields consistent performance gains and achieves SOTA results, suggesting that structured, Git-style memory is a key ingredient for effective long-horizon reasoning in autonomous agents.